OpenStudy (loser66):

After calculating, I got this rI -A is \[\left[\begin{matrix}3&-1&-2\\1&2&1\\2&1&3\end{matrix}\right]\] I forgot how to find eigenvector from that, help me, please

OpenStudy (loser66):

@e.mccormick

OpenStudy (loser66):

I replace eigenvalue into the RI -A to get the matrix above. And ...I need you help, hehehe

OpenStudy (e.mccormick):

RI? Mean \(\lambda I \)?

OpenStudy (loser66):

yup

OpenStudy (e.mccormick):

OK. Did you get the determinant?

OpenStudy (loser66):

why? do we have to have it?

OpenStudy (e.mccormick):

That gets you the eigenvalues.

OpenStudy (loser66):

ok, let me show you the original A \[\left[\begin{matrix}1&1&2\\1&2&1\\2&1&1\end{matrix}\right]\]

OpenStudy (loser66):

I have characteristic equation is \[\lambda^3-4\lambda^2-\lambda+4=0\\(\lambda+1)(\lambda -4)(\lambda -1)=0\] so the eigenvalues are -1, 1 and 4 the matrix above is got from eigenvalue 4 and stuck

OpenStudy (e.mccormick):

OK, so you did the value part already.

OpenStudy (loser66):

yes

OpenStudy (e.mccormick):

For each eigenvalue, you solve \((A-\lambda I)=0\) by Gaussian Elimination.

OpenStudy (loser66):

rref?

OpenStudy (e.mccormick):

missed an \(\vec{x}\) after the )....

OpenStudy (loser66):

ok, we know it, doesn't matter, go ahead

OpenStudy (e.mccormick):

So for the first value, -1, we have: \(\begin{bmatrix} 1-\lambda & 1 & 2\\ 1 & 2-\lambda & 1\\ 2 & 1 & 1-\lambda \end{bmatrix}\vec{x}=0 \implies \) \(\begin{bmatrix} 1+1 & 1 & 2\\ 1 & 2+1 & 1\\ 2 & 1 & 1+1 \end{bmatrix}\vec{x}=0 \implies \) \(\left[ \begin{array}{ccc|c} 2 & 1 & 2&0\\ 1 & 3 & 1&0\\ 2 & 1 & 2&0\\ \end{array} \right]\)

OpenStudy (loser66):

why? I ask for 4 , not for -1 why do you force me to do the whole set??

OpenStudy (e.mccormick):

And you put that in ref, not rref. Upper triangular form.

OpenStudy (e.mccormick):

I was just picking it as an example. If you want to do another, do so.

OpenStudy (loser66):

I want 4. you know why? because it 's hardest.

OpenStudy (loser66):

If I can solve the hardest one. the easier one will be "no problem" , right?

OpenStudy (e.mccormick):

I don't know that any of them are really any harder or easier. At that point, it is basically putting it into parametric form.

OpenStudy (loser66):

hey, what is parametric form?

OpenStudy (e.mccormick):

Know parametic equations? Where you represent all variables as combinations of one of them?

OpenStudy (e.mccormick):

Here is an explantion of what I mean: http://www.scss.tcd.ie/Rozenn.Dahyot/CS1BA1/SolutionEigen.pdf The lower part of this shows it.

OpenStudy (loser66):

yes, know it from cal3 but don't see the link to linear algebra

OpenStudy (loser66):

Ok, I think I am ok now. Thanks for the link @e.mccormick. I will show you the answer. hihi

OpenStudy (loser66):

one more thing, my prof wants me apply RI -A , not A - RI . ha!! each prof gives out his way to apply. I have to go back and forth between the 2

OpenStudy (e.mccormick):

They are the same in the end. It is just if you start one way, you use that way all through.

OpenStudy (loser66):

hey, do we always have a row =0 from ref?

OpenStudy (loser66):

at least 1 , at most 2 if it is 3x3 matrix?

OpenStudy (e.mccormick):

It is implied if you leave it off. The main thing to realize is that you are solving an Ax=b type equation where b is the zero vector and you want to get at x.

OpenStudy (loser66):

you confused me, because if it's so, x = A^-1 b =0

OpenStudy (e.mccormick):

When the x is a series of variables, it distributes across the matrix. \(\begin{align*} ax_1 + bx_2 + cx_3 &= 0 \\ dx_1 + ex_2 + fx_3 &= 0 \\ gx_1 + hx_2 + ix_3 &= 0 \end{align*}\) becomes: \(\begin{bmatrix} a & b& c \\ d&e&f\\ g&h&i \end{bmatrix} \left \langle \begin{matrix} x_1\\x_2\\x_3 \end{matrix} \right \rangle=0\)

OpenStudy (e.mccormick):

Or more accurately, \(=\vec{0}\) on the right of that....

OpenStudy (loser66):

but cannot pick \(\vec 0\)

OpenStudy (e.mccormick):

The 0 vector on the right is the same thing as having the augmented matrix with zeros. All the zeros do is let you put say \(x_1\) in terms of \(x_2\) and \(x_3\).

OpenStudy (loser66):

got you. thanks e.mccormick